OpenStudy (anonymous):

3/8 + 1/12 What is one denominator you could use to add the following two fractions? Find the sum of the two fractions. Show work please

OpenStudy (texaschic101):

if you have trouble finding a common denominator, you can just multiply the 2 denominators. 8 x 12 = 96....in some cases, it may not be the smallest common denominator, but at the end of the problem, you can reduce. The smallest would be 24...because 8 goes into 24 and 12 goes into 24

OpenStudy (anonymous):

in order to add the fractions, they need to have the same denominator, LCD. 8 and 12 share a factor of 24. so multiply 3/8*4/4 and 1/12*24/24

OpenStudy (anonymous):

1/12*2/2

OpenStudy (anonymous):

sooooooo? the answer is 8 and 12?

OpenStudy (anonymous):

then add x/24+x/24

OpenStudy (anonymous):



OpenStudy (anonymous):

thx guys ill be putting up another question soon

OpenStudy (texaschic101):

3/8 + 1/12 --- common denominator is 24 basically you are looking for an equivalent fraction with a denominator of 24. 3/8 --- what do you need to multiply 8 by to get 24 ? That would be 3. SO multiply numerator and denominator by 3. 3/8 * 3/3 = 9/24 same with 1/12...what do you have to multiply 12 by to get 24 ? That would be 2, so multiply numerator and denominator by 2. 1/12 * 2/2 = 2/24 now we have : 9/24 + 2/24 = 11/24 Little tip...equivalent fractions can be found by multiplying numerator and denominator by the same number.